# Spectral Properties of Irradiated Circumbinary Disks around Binary Black Holes Governed by Hydrogen Opacities Dependent on Temperature and Density

### v1

@@ Abstract
We study the thermal and spectral properties of irradiated circumbinary disks (CBDs) around binary black holes (BBHs), using analytic, hydrogen-based opacity models that capture key dependencies on temperature, density, and ionization. We solve the vertically hydrostatic energy balance equations with Rosseland mean opacities from free-free absorption, bound-free absorption, and electron scattering processes, combined with ionization fractions derived by the Saha equation. Four opacity models are considered, including a reference model with no physical opacity, constructed by Lee et al. (2024), and three physically motivated alternatives. The midplane temperature profiles show significant variation across models, while the surface temperature remains largely unchanged in regions dominated by viscous heating. Opacity effects become pronounced in the outer disk, where irradiation reprocessing shapes the IR-optical continuum. Inclusion of bound-free opacity introduces a noticeable flattening and a mid-frequency peak in the spectral energy distribution. We also compute spectra of a triple disk system including the CBD and two accreting minidisks. The high-frequency peak arises from the hot minidisks, while the low-frequency excess originates from irradiated outer CBD layers. By comparing model spectra with detection limits of Subaru, JWST, and Swift, we find that systems within $\sim 10$ Mpc may show detectable IR excess from the CBD. Our results highlight the need for accurate opacity modeling to interpret electromagnetic signatures of black hole mergers and support future integration of opacity tables with metallicity.
@@
1 Introduction
Stellar-mass binary black holes (BBHs), composed of two black holes with masses in the range ${3\,M_{\odot}}\leq M\leq{100\,M_{\odot}}$, can form either through the isolated binary evolution of massive stars (Marchant et al., 2016) or through dynamical interactions in dense stellar systems such as globular clusters (Morscher et al., 2015). Since the first direct detection of a black hole merger by the LIGO Observatory (event GW150914) (Abbott et al., 2016), a growing number of such events have been reported (Abbott et al., 2023; Nitz et al., 2023), including those suggesting the existence of intermediate-mass black holes (IMBHs). On a larger scale, supermassive black hole (SMBH) binaries are thought to reside at the centers of active galactic nuclei (AGNs) and may play a key role in galaxy evolution as well as structure formation in the early universe. At all these mass scales, the presence of circumbinary disks (CBDs) around BBHs provides an important way to probe binary evolution via electromagnetic (EM) signals complementary to gravitational waves (GWs).
The thermal and spectral properties of a CBD depend critically on its internal heating and radiative cooling processes. In particular, CBDs are expected to be heated by viscous dissipation and, in some cases, by irradiation from an accreting minidisk around each black hole. A recent study by Lee et al. (2024) investigated the effect of irradiation from such minidisks in shaping the CBD temperature profile and spectrum, showing that reprocessed emission can lead to a characteristic infrared (IR) peak in the spectral energy distribution (SED). However, their analysis assumed the disk surface temperature equals to the disk midplane temperature, resulting in a fixed optical depth $\tau\lesssim 1$, thereby neglecting the strong temperature and ionization dependence of opacity in cooler, denser regions of the CBD.
@@ 1 Introduction
$Q_{\rm vis}+Q_{\rm irr}=Q_{\rm rad},$ (1)
where $Q_{\rm vis}$ is the viscous heating rate, $Q_{\rm irr}$ is the irradiation heating rate, and $Q_{\rm rad}$ is the radiative cooling rate depending on $\kappa(\rho,T_{\rm c},x_{e}(T_{\rm c}))$. The resulting models allow us to quantify how microphysical opacity prescriptions shape the thermal state and spectral signature of the CBD with and without irradiation.
@@
In Section 2, we present the basic framework for modeling a one-dimensional (1D), geometrically thin, optically thick, irradiated CBD with the detailed description of the three opacity prescriptions. In Section 3, we numerically solve the energy equation of the CBD. Among the three opacity models, the model involving combined free-free absorption, bound-free absorption, and electron scattering processes requires numerical integration due to their nonlinear dependence on the density, temperature, and ionization degree. In Section 4, we present the resulting radial profiles of disk temperature, the emergent spectra from the CBD, and the combined spectra from the CBD and minidisks, referred to as the triple disk spectra. Sections 5 and 6 are devoted to discussion and conclusions, respectively.
@@
2 Model
We consider a 1D, axisymmetric, and steady-state CBD surrounding a BBH system. The CBD is assumed to be optically thick and geometrically thin, and to evolve under viscous dissipation in the presence of external irradiation from the minidisks.
@@ 2 Model
The total mass of the binary is defined as $M=M_{1}+M_{2}$, where $M_{1}$ and $M_{2}$ are the masses of the primary and secondary black holes, respectively. The binary follows a circular orbit with a semi-major axis $a=a_{0}\,r_{\rm S}$, where $a_{0}$ is a dimensionless scaling factor and the Schwarzschild radius, $r_{S}$, is defined as
@@
with $1.6\lesssim C_{\rm gap}\lesssim 4$ (Artymowicz & Lubow, 1994; Dittmann et al., 2023). We adopt $C_{\rm gap}=2$ as a fiducial value throughout this paper.
@@
is the optical depth, with $\Sigma$ being the surface density of the disk. In our formulation, the opacity $\kappa(\rho,T_{\rm c},x_{e})$ consists of a sum of contributions from electron scattering, free-free absorption, and bound-free absorption (see Section 3 for details). Equating equation (4) with equation (5) yields the relation between the surface and midplane temperatures:
@@
with the viscosity parameter $\alpha_{\rm SS}$, where $c_{\rm s}$ is the sound speed. Assuming the CBD is in hydrostatic equilibrium for the direction perpendicular to the disk plane, the sound speed $c_{\rm s}$ is given by
@@
$\frac{\tau_{\rm vis}\left(r_{\rm in}\right)}{P_{\rm orb}}$ $\sim$ $2.0\times 10^{4}\left(\frac{C_{\rm gap}}{2}\right)^{3}\left(\frac{a_{0}}{1000}\right)^{3/2}\left(\frac{\alpha_{\rm SS}}{0.1}\right)^{3}\left(\frac{H/r}{0.01}\right)^{2},$ (13)
@@
Since equation (13) suggests that $\tau_{\rm vis}\gg P_{\rm orb}$, we assume that the binary system with triple disk composed of CBD, CPD, and CSD is in a quasi-steady state. This allows to impose on the following relation on the accretion rates between CBD, CPD, and CSD:
@@
2.1 Disk Opacity
@@ 3 Method
3.1 Optical depth prescriptions
Considering that the CBD is modeled by the standard disk model, the surface density is given by
$\Sigma$ $=$ $\frac{\dot{M}}{3\pi\nu}=\sqrt{2}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{\sqrt{C_{\rm gap}a_{0}}}\frac{1}{\kappa_{\rm es}}\frac{1}{Y^{2}\xi^{1/2}},$ (29)
where we used equations (3), (9), (11), (12) and (27) for the derivation. The mass density is then given by
$\rho=\frac{1}{2}\frac{\Sigma}{H}$ $=$ $\frac{1}{2}\frac{r}{H}\frac{\Sigma}{r_{\rm S}}\frac{r_{\rm S}}{r_{\rm in}}\frac{r_{\rm in}}{r}=\frac{1}{2}\frac{1}{Y\xi}\frac{\Sigma}{r_{\rm S}}\frac{1}{C_{\rm gap}a_{0}}=\frac{1}{\sqrt{2}}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{(C_{\rm gap}a_{0})^{3/2}}\frac{1}{r_{\rm S}\kappa_{\rm es}}\frac{1}{Y^{3}\xi^{3/2}}.$ (30)
@@ 3.1 Optical depth prescriptions
$\kappa_{\rm a}$ $=$ $\frac{1}{\sqrt{2}}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{(C_{\rm gap}a_{0})^{3/2}}\frac{\kappa_{\rm ff,0}}{\kappa_{\rm es}}\frac{1}{r_{\rm S}}x_{e}^{2}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/2}\left[1+\frac{1-x_{e}}{x_{e}^{2}}\frac{\kappa_{\rm bf,0}}{\kappa_{\rm ff,0}}\right]\frac{\xi^{2}}{Y^{10}}.$ (33)
@@
$\kappa_{2}$ $=$ $\frac{1}{2^{1/4}}\left(\frac{\dot{m}}{\alpha_{\rm SS}}\right)^{1/2}\frac{1}{(C_{\rm gap}a_{0})^{3/4}}\left(\frac{\kappa_{\rm ff,0}}{r_{\rm S}}\right)^{1/2}x_{e}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/4}\frac{\xi}{Y^{5}},$ (34)
$\kappa_{3}$ $=$ $\frac{1}{\sqrt{2}}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{(C_{\rm gap}a_{0})^{3/2}}\frac{\kappa_{\rm ff,0}}{\kappa_{\rm es}}\frac{1}{r_{\rm S}}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/2}\left[x_{e}^{2}+(1-x_{e})\frac{\kappa_{\rm bf,0}}{\kappa_{\rm ff,0}}\right]\frac{\xi^{2}}{Y^{10}}$ (35)
$\times$ $\Biggr{[}1+\sqrt{2}\frac{\alpha_{\rm SS}}{\dot{m}}(C_{\rm gap}a_{0})^{3/2}r_{\rm S}\frac{\kappa_{\rm es}^{2}}{\kappa_{\rm ff,0}}\left(\frac{T_{0}}{1+x_{e}}\right)^{7/2}\left[\frac{1}{x_{e}^{2}+(1-x_{e})(\kappa_{\rm bf,0}/\kappa_{\rm ff,0})}\right]\frac{Y^{10}}{\xi^{2}}\Biggr{]}^{1/2}.$
@@
$\tau_{30}$ $=$ $\frac{1}{2}\left(\frac{\dot{m}}{\alpha_{\rm SS}}\right)^{2}\frac{1}{(C_{\rm gap}a_{0})^{2}}\frac{\kappa_{\rm ff,0}}{\kappa_{\rm es}^{2}}\frac{1}{r_{\rm S}}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/2}\left[x_{e}^{2}+(1-x_{e})\frac{\kappa_{\rm bf,0}}{\kappa_{\rm ff,0}}\right],$
$\tau_{31}$ $=$ $\sqrt{2}\frac{\alpha_{\rm SS}}{\dot{m}}(C_{\rm gap}a_{0})^{3/2}\frac{\kappa_{\rm es}^{2}r_{\rm S}}{\kappa_{\rm ff,0}}\left(\frac{T_{0}}{1+x_{e}}\right)^{7/2}\left[\frac{1}{x_{e}^{2}+(1-x_{e})(\kappa_{\rm bf,0}/\kappa_{\rm ff,0})}\right].$
3.2 Energy equation

### v2

@@ Abstract
We study the thermal and spectral properties of irradiated circumbinary disks (CBDs) around binary black holes (BBHs), using analytic, hydrogen-based opacity models that capture dependencies on temperature, density, and ionization. We solve the vertical hydrostatic equilibrium and energy balance, assuming gas pressure only, using Rosseland-mean opacities from free–free and bound–free absorption plus electron scattering, with ionization fractions given by the Saha equation. Four opacity models are considered, including a reference model with no physical opacity, constructed by Lee et al. (2024), and three physically motivated alternatives. The midplane temperature profiles show significant variation across models, while the surface temperature remains largely unchanged in regions dominated by viscous heating. Opacity effects become pronounced in the outer disk, where irradiation reprocessing shapes the IR-optical continuum. Bound-free opacity introduces flattening and a mid-frequency peak in the spectral energy distribution. We compute spectra of a triple disk system including the CBD and two accreting minidisks. The high-frequency peak arises from the hot minidisks, while the low-frequency excess originates from irradiated outer CBD layers. Comparing model spectra with detection limits of Subaru, JWST, and Swift, we find that BBH systems within $\sim 10$ Mpc can exhibit a detectable IR excess. Our results highlight the need for physically consistent opacity modeling to interpret electromagnetic (EM) signatures of BBHs approaching coalescence and support integration of metallicity-dependent opacity tables. Our opacity-informed framework for irradiated CBDs provides an EM template for identifying stellar- to intermediate-mass BBHs in a mass range sparsely sampled by LISA, thereby bridging the gravitational-wave–EM gap with testable IR/optical signatures.
@@
I Introduction
Stellar-mass binary black holes (BBHs), composed of two black holes with individual masses between $3\,M_{\odot}$ and $100\,M_{\odot}$, can form either through the isolated binary evolution of massive stars (Marchant et al., 2016) or through dynamical interactions in dense stellar systems such as globular clusters (Morscher et al., 2015). Since the first direct detection of a black hole merger by the LIGO Observatory (event GW150914) (Abbott et al., 2016), a growing number of such events have been reported (Abbott et al., 2023; Nitz et al., 2023), including those suggesting the existence of intermediate-mass black holes (IMBHs). On a larger scale, supermassive black hole (SMBH) binaries are thought to reside at the centers of active galactic nuclei (AGNs) and may play a key role in galaxy evolution as well as structure formation in the early universe. At all these mass scales, the presence of circumbinary disks (CBDs) around black-hole binaries provides an important way to probe binary evolution via electromagnetic (EM) emission complementary to gravitational waves (GWs).
The thermal and spectral properties of a CBD depend critically on its internal heating and radiative cooling processes. In particular, CBDs are expected to be heated by viscous dissipation and, in some cases, by irradiation from an accreting minidisk around each black hole. A physically consistent treatment of irradiation is therefore essential for predicting CBD spectra and EM observables.
In this context, the single–black-hole disk literature provides well-tested baselines for irradiation physics. Early studies established the framework of X-ray–heated, reprocessing disks (Hayakawa, 1981), and general-relativistic “returning radiation” can self-irradiate the outer disk around a Schwarzschild black hole (Sanbuichi et al., 1993), building on earlier Newtonian treatments of flared, self-irradiated disks (Fukue, 1992). In stellar-mass X-ray binaries, these ideas matured into quantitative reprocessing models that link optical/infrared (IR) light to the X-ray luminosity and help explain outburst decays and the maintenance of the hot state (van Paradijs & McClintock, 1994; King & Ritter, 1998; Dubus, 1999; Dubus et al., 2001). As contemporary observational anchors, irradiated disk models have been fit jointly to X-ray and optical data of ultraluminous X-ray sources (e.g. Sutton et al. 2014), and recent multi-wavelength campaigns in single-black-hole systems now recover ultraviolet (UV)/optical lags (Oknyansky et al., 2025) and spectral energy distributions (SEDs) consistent with X-ray thermal reprocessing (Yoshitake et al., 2024). These single-black-hole baselines motivate how we incorporate irradiation from the minidisks in the CBD context below.
A recent study by Lee et al. (2024) investigated the effect of irradiation from such minidisks in shaping the CBD temperature profile and spectrum, showing that reprocessed emission can lead to a characteristic IR peak in the SED. However, their analysis assumed the disk surface temperature equals the disk midplane temperature, equivalent to adopting a fixed effective optical depth $\tau=16/3$, thereby neglecting the strong temperature and ionization dependence of opacity in cooler, denser regions of the CBD.
@@ I Introduction
In BBH systems, ionizing UV photons and X-rays emitted by the minidisks can strongly modify the ionization state of the CBD surface layers, and hence the relevant opacities. In the irradiation-dominated surface layers of a CBD, such hard photons can suppress dust and reduce the importance of metal-line absorption by driving the gas toward high ionization (Barvainis, 1987). In the irradiation-dominated layers relevant for CBD reprocessing, the opacity can therefore be treated as effectively hydrogen-dominated, together with electron scattering. Motivated by this, and to isolate the irradiation/ionization physics relevant to CBD reprocessing, we adopt a clean baseline model with a pure-hydrogen composition: we use analytic Kramers-type prescriptions for $\kappa_{\rm ff}$ and $\kappa_{\rm bf}$ together with $\kappa_{\rm es}$, and compute the ionization fraction self-consistently from the Saha equation. This baseline allows us to quantify how the opacity prescription controls (i) the midplane–surface temperature contrast and (ii) the resulting IR/optical emission from the outer CBD. We revisit the role of metals, dust, and tabulated opacities in Section V.1.
@@
$Q_{\rm vis}+Q_{\rm irr}+Q_{\rm tid}=Q_{\rm rad},$ (1)
where $Q_{\rm vis}$ is the viscous heating rate, $Q_{\rm irr}$ is the irradiation heating rate, $Q_{\rm tid}$ is the heating rate due to dissipation of tidally excited density waves near the cavity rim, and $Q_{\rm rad}$ is the radiative cooling rate depending on $\kappa(\rho,T_{\rm c},x_{e}(T_{\rm c}))$.
Throughout this paper, we adopt a one-dimensional (1D), azimuthally and orbit-averaged steady framework [^1] This time-averaged treatment is justified because the viscous timescale at the inner edge, $\tau_{\rm vis}(r_{\rm in})$, is much longer than the binary orbital period $P_{\rm orb}$, as quantified in equation (13).. In all CBD models presented in this paper we set $Q_{\rm tid}=0$. Analytic studies show that the dissipation of these waves is confined to a narrow annulus of width $\sim$ a few scale heights $H$ at $r\simeq r_{\rm in}$ (Lodato et al., 2009; Rafikov, 2013). For the geometrically thin disks considered here, with $H/r\lesssim 10^{-2}$, this tidally heated region occupies only a small fraction $\sim H/r\ll 1$ of the CBD surface and thus contributes negligibly to the disk-wide energy budget and to the outer-disk surface temperature that sets the global SED. This choice is also consistent with our surface-density prescription, which ignores tidal torques away from the rim (see equations (29) and (30) of Section III.1 for details). The resulting models allow us to quantify how microphysical opacity prescriptions shape the thermal state and spectral signature of the CBD with and without irradiation.
@@
In Section II, we present the basic framework for modeling a 1D, geometrically thin, optically thick, irradiated CBD with the detailed description of the three opacity prescriptions. In Section III, we numerically solve the energy equation of the CBD. Among the three opacity models, the model involving combined free-free absorption, bound-free absorption, and electron scattering processes requires numerical integration due to their nonlinear dependence on the density, temperature, and ionization degree. In Section IV, we present the resulting radial profiles of disk temperature, the emergent spectra from the CBD, and the combined spectra from the CBD and minidisks, referred to as the triple disk spectra. Sections V and VI are devoted to discussion and conclusions, respectively.
@@
II Model
We consider a 1D, axisymmetric, and steady-state CBD surrounding a BBH system, computed from azimuthally and orbit-averaged equations that neglect explicitly time-dependent and non-axisymmetric components of the binary tidal potential. The CBD is assumed to be optically thick and geometrically thin, and to evolve under viscous dissipation in the presence of external irradiation from the minidisks. In this framework, the binary affects the CBD structure only through the geometric truncation at the cavity edge described by equation (3) below and through the irradiation source term defined by equation (LABEL:eq:qirr) in Section III.2
@@ II Model
The total mass of the binary is defined as $M=M_{1}+M_{2}$, where $M_{1}$ and $M_{2}$ are the masses of the primary and secondary black holes, respectively. The binary follows a circular orbit with a semi-major axis $a=a_{0}\,r_{\rm S}$, where $a_{0}$ is a dimensionless scaling factor and the Schwarzschild radius, $r_{\rm S}$, is defined as
@@
with $1.6\lesssim C_{\rm gap}\lesssim 4$ (Artymowicz & Lubow, 1994; Dittmann et al., 2023). We adopt $C_{\rm gap}=2$ as a fiducial value throughout this paper. The binary enters explicitly through the geometric truncation of the CBD at the cavity edge, $r_{\rm in}$. This prescription suppresses local viscous heating for $r<r_{\rm in}$ and limits the CBD’s emitted area and the angles at which the CBD is irradiated. Apart from this truncation radius, our models do not consider time-dependent, non-axisymmetric tidal heating or inner torque terms.
@@
is the optical depth, with $\Sigma$ being the surface density of the disk. In our formulation, the opacity $\kappa(\rho,T_{\rm c},x_{e})$ consists of a sum of contributions from electron scattering, free-free absorption, and bound-free absorption (see Section III for details). Equating equation (4) with equation (5) yields the relation between the surface and midplane temperatures:
@@
with the viscosity parameter $\alpha_{\rm SS}$, where $c_{\rm s}$ is the sound speed. Assuming that the CBD is vertically supported by gas pressure and is in hydrostatic equilibrium for the direction perpendicular to the disk plane, the sound speed $c_{\rm s}$ is given by
@@
$\frac{\tau_{\rm vis}\left(r_{\rm in}\right)}{P_{\rm orb}}$ $\sim$ $2.0\times 10^{4}\left(\frac{C_{\rm gap}}{2}\right)^{3}\left(\frac{a_{0}}{1000}\right)^{3/2}\left(\frac{\alpha_{\rm SS}}{0.1}\right)^{3}$ (13)
$\times$ $\left(\frac{H/r}{0.01}\right)^{2},$
@@
Since equation (13) suggests that $\tau_{\rm vis}\gg P_{\rm orb}$, we assume that the binary system with triple disk composed of CBD, CPD, and CSD is in a quasi-steady state. This allows us to impose on the following relation on the accretion rates between CBD, CPD, and CSD:
@@
In the equal-mass limit $q=1$, symmetry implies that the instantaneous accretion rates onto the two minidisks are comparable, $\dot{M}_{2}/\dot{M}_{1}\simeq q$, as found in hydrodynamic simulations of circumbinary accretion (Farris et al., 2014; Duffell et al., 2020). Because our study focuses on this limit and is motivated by these findings, we adopt the prescription of $\dot{M}_{2}/\dot{M}_{1}=q$ throughout this paper.
@@
Throughout this work, the vertical support and sound speed are computed with an ideal gas equation of state, with the mean molecular weight determined by the Saha ionization fraction. Consequently, the pressure that enters the hydrostatic balance and the energy equation is gas pressure alone. Neither the vertical force balance nor the equation of state includes radiation pressure, and magnetic pressure is likewise neglected. This treatment is consistent with our focus on cold, geometrically thin, optically thick CBDs, where $T_{\rm c}\sim 10^{3}$-$10^{5}\,{\rm K}$, and irradiation primarily affects the surface layers rather than the midplane structure. We discuss the applicability and limitations of this approximation in Section V.
II.1 Disk Opacity
@@ III Method
III.1 Optical depth prescriptions
Following Rafikov (2013), the conservation law of the radial angular momentum flux $F_{J}$ in a steady state is given by $F_{J}=3\pi\nu\Sigma{l}-\dot{M}l$, where $l=r^{2}\Omega$ and a constant inner-boundary contribution $F_{J}$ that encapsulates the binary torque. We then obtain the surface density profile
$\Sigma=\Sigma_{0}\Bigl[1+\epsilon_{\rm bin}\xi^{-1/2}\Bigr],$ (29)
where $\epsilon_{\rm bin}=F_{J}/(\dot{M}\sqrt{GMr_{\rm in}})$ is the binary-torque parameter, and $\Sigma_{0}$ is the surface density of the standard disk model:
$\Sigma_{0}$ $=$ $\frac{\dot{M}}{3\pi\nu}=\sqrt{2}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{\sqrt{C_{\rm gap}a_{0}}}\frac{1}{\kappa_{\rm es}}\frac{1}{Y^{2}\xi^{1/2}},$ (30)
Here, equations (3), (9), (11), (12), and (27) are used for the derivation. Equation (30) makes explicit how a nonzero inner angular momentum flux inherited from the binary ($\epsilon_{\rm bin}\neq 0$) produces an inner pile-up ($\epsilon_{\rm bin}>0$) or depletion ($\epsilon_{\rm bin}<0$) [^2] In the circumbinary systems of interest, the binary is expected to lose angular momentum to the CBD, so that the corresponding inner torque and $F_{J}$ are non-negative ($\epsilon_{\rm bin}\geq 0$); we therefore do not consider the depletion regime with $\epsilon_{\rm bin}<0$ in this work., while reducing to $\Sigma_{0}$ at the zero-torque limit ($\epsilon_{\rm bin}=0$). For simplicity, we take $\epsilon_{\rm bin}=0$ throughout this paper.
The mass density is then given by
$\rho=\frac{1}{2}\frac{\Sigma}{H}$ $=$ $\frac{1}{2}\frac{r}{H}\frac{\Sigma}{r_{\rm S}}\frac{r_{\rm S}}{r_{\rm in}}\frac{r_{\rm in}}{r}=\frac{1}{2}\frac{1}{Y\xi}\frac{\Sigma}{r_{\rm S}}\frac{1}{C_{\rm gap}a_{0}}\$(31)
$=$ $\frac{1}{\sqrt{2}}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{(C_{\rm gap}a_{0})^{3/2}}\frac{1}{r_{\rm S}\kappa_{\rm es}}\frac{1}{Y^{3}\xi^{3/2}}.$
@@ III.1 Optical depth prescriptions
$\kappa_{\rm a}$ $=$ $\frac{1}{\sqrt{2}}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{(C_{\rm gap}a_{0})^{3/2}}\frac{\kappa_{\rm ff,0}}{\kappa_{\rm es}}\frac{1}{r_{\rm S}}$ (34)
$\times$ $x_{e}^{2}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/2}\left[1+\frac{1-x_{e}}{x_{e}^{2}}\frac{\kappa_{\rm bf,0}}{\kappa_{\rm ff,0}}\right]\frac{\xi^{2}}{Y^{10}}.$
@@
$\kappa_{2}$ $=$ $\frac{1}{2^{1/4}}\left(\frac{\dot{m}}{\alpha_{\rm SS}}\right)^{1/2}\frac{1}{(C_{\rm gap}a_{0})^{3/4}}\left(\frac{\kappa_{\rm ff,0}}{r_{\rm S}}\right)^{1/2}$ (35)
$\times$ $x_{e}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/4}\frac{\xi}{Y^{5}},$
$\kappa_{3}$ $=$ $\frac{1}{\sqrt{2}}\frac{\dot{m}}{\alpha_{\rm SS}}\frac{1}{(C_{\rm gap}a_{0})^{3/2}}\frac{\kappa_{\rm ff,0}}{\kappa_{\rm es}}\frac{1}{r_{\rm S}}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/2}$ (36)
$\times$ $\left[x_{e}^{2}+(1-x_{e})\frac{\kappa_{\rm bf,0}}{\kappa_{\rm ff,0}}\right]\frac{\xi^{2}}{Y^{10}}$
$\times$ $\Biggr[1+\sqrt{2}\frac{\alpha_{\rm SS}}{\dot{m}}(C_{\rm gap}a_{0})^{3/2}r_{\rm S}\frac{\kappa_{\rm es}^{2}}{\kappa_{\rm ff,0}}\left(\frac{T_{0}}{1+x_{e}}\right)^{7/2}$
$\times$ $\left[\frac{1}{x_{e}^{2}+(1-x_{e})(\kappa_{\rm bf,0}/\kappa_{\rm ff,0})}\right]\frac{Y^{10}}{\xi^{2}}\Biggr]^{1/2}.$
@@
$\tau_{30}$ $=$ $\frac{1}{2}\left(\frac{\dot{m}}{\alpha_{\rm SS}}\right)^{2}\frac{1}{(C_{\rm gap}a_{0})^{2}}\frac{\kappa_{\rm ff,0}}{\kappa_{\rm es}^{2}}\frac{1}{r_{\rm S}}\left(\frac{1+x_{e}}{T_{0}}\right)^{7/2}$
$\times$ $\left[x_{e}^{2}+(1-x_{e})\frac{\kappa_{\rm bf,0}}{\kappa_{\rm ff,0}}\right],$
$\tau_{31}$ $=$ $\sqrt{2}\frac{\alpha_{\rm SS}}{\dot{m}}(C_{\rm gap}a_{0})^{3/2}\frac{\kappa_{\rm es}^{2}r_{\rm S}}{\kappa_{\rm ff,0}}\left(\frac{T_{0}}{1+x_{e}}\right)^{7/2}$
$\times$ $\left[\frac{1}{x_{e}^{2}+(1-x_{e})(\kappa_{\rm bf,0}/\kappa_{\rm ff,0})}\right].$
At fixed $(\xi,Y,M)$, equations (LABEL:eq:kff2)–(34) show that the absorption opacities increase as the binary semi-major axis $a\,(=a_{0}\,r_{\rm S})$ becomes larger. This happens because the dynamical temperature scale decreases as $T_{0}\propto a^{-1}$ and the disk density scales as $\rho\propto a^{-3/2}$. When these scalings are inserted into the Kramers-type Rosseland means, one finds that $\kappa_{\rm ff}$ and $\kappa_{\rm bf}$ grow approximately as $\kappa_{\rm ff},\kappa_{\rm bf}\propto a^{2}$, apart from the ionization factors $x_{e}$ and $1-x_{e}$.
As $a/r_{\rm S}$ increases, the ionization fraction $x_{e}$ obtained from equation (19) decreases. As a result, $\kappa_{\rm ff}$ reaches a broad maximum at intermediate $a/r_{\rm S}$ and then declines, while $\kappa_{\rm bf}$ continues to rise and eventually dominates the total opacity. Figure 2 illustrates this behavior. At small separations, electron scattering represented by $x_{\rm e}\kappa_{\rm es}$ provides the largest contribution to the total opacity. As the binary separation increases toward $a/r_{\rm S}\sim 10^{2}$, the free–free opacity $\kappa_{\rm ff}$ becomes comparable to $x_{e}\kappa_{\rm es}$ and both contribute significantly. For even larger separations, the bound–free opacity $\kappa_{\rm bf}$ overtakes these components and dominates $\kappa_{\rm tot}$, which then increases steeply. This is the regime relevant for our models. Thus, the dominant source of opacity in the CBD changes systematically with the binary separation.
Figure 2: Analytic pure-hydrogen Rosseland-mean opacity $\kappa$ contributions as a function of the semi-major axis normalized by the Schwarzschild radius, $a/r_{\rm S}$. The opacities are evaluated for a pure-hydrogen CBD with $M=100\,M_{\odot}$, $Y=0.01$, and $\xi=10$. Colors denote the contributions from electron scattering multiplied by the ionization fraction ($x_{e}\kappa_{\rm es}$, black), free-free absorption ($\kappa_{\rm ff}$, red), and bound-free absorption ($\kappa_{\rm bf}$, blue). The black dashed curve shows the total opacity, $\kappa=x_{e}\kappa_{\rm es}+\kappa_{\rm ff}+\kappa_{\rm bf}$
III.2 Energy equation